Felix Jones, a recent engineering graduate, expects a starting salary of $35,000 per year. His future employer has averaged 5% per year in salary in- creases for the last several years. If inflation is esti- mated to be 4% per year for the next 3 years, how much, in year-1 dollars, will Felix be earning each year? What is the differential inflation rate in Felix’s salary?
Answer #1

Year

Expected salary Present value factor based on inflation Value in Year 1 dollars Average inflation
0 35,000 1

35,000

-
1 36,750 0.962 35,336.54 0.009615
2 38,587.50 0.925 35,676.31 0.009615
3 40,516.88 0.889 36,019.35 0.009615

Currently Felix's salary is $35,000 i.e. salary in Year 0

in year 1, expected salary = 35,000* (1 + 0.05) = $36, 750 present value factor = 1 +0,04 = 0.962 value in year 1 dollars = 3

in year 2 expected salary = 35,000 * (1 +0.05)= $38,587.50 present value factor = (2004)2 = 0.925 value in year 1 dollars = 3

in year3 expected salary = 35,000 * (1 +0.05)3 = $40,516.88 present value factor = (1 +004)3 = 0.889 value in year 1 dollars

Average inflation rate is 0.9615%